600 mL of D5W was to infuse in 5 hours at 33 macrott/min. The drop factor is 20 gtt/mL. After 2 hours, your notice 350 mL remaining. Recalculate rate ____ and the percentage of change and course of action
Answer #1

The drop factor 20 gtt/ml indicates the infusion is on a macro drip.

Intra Venous (IV) drip rate is calculated using the formula below

Total Volume (ml)       = ml per hour

Total time (in hours)

350 ml    = 116.66 ml/hr

3 hours

The IV flow rate in drops per minute (gtt/min)

Volume (ml)        \times drop factor (gtt/min) = IV flow rate

Time (min)

350 ml           \times 20 = 38.88 = 39 gtt/min

180 minutes

The infusion should be administered for 39 drops per minute to infuse 350 ml of D5W in the next 3 hours.

Increase= New number- original number

Percentage of Increase = Increase divided by original number multiplied with 100. That is 6/33*100 = 18.18%